Formula for the force of attraction between two parallelmoving protons

Grade:12th pass

1 Answers

Rajdeep
5 years ago
HEY THERE!
 
Here’s the general formula:
 
\small F = k_{e}\frac{q_{1}q_{2}}{r^{2}}\\\\ k_{e} = 8.99 \times 10^{9} Nm^{2}C^{-2}
 
Where, \small q_{1}, q_{2} are the charges of proton (both are same, = +1.6 x 10-19 C) and r is the distance between the two parallel moving protons.
 
So, we have :
 
\small F = 8.99\times 10^{9} \times \frac{(1.6\times 10^{-19})^{2}}{r^{2}} \\\\= \frac{2.301\times10^{-28}}{r^{2}}
 
THANKS!